OpenStudy (anonymous):

the radius of a circle measures 20 centimeters. what is the length of its diameter?

OpenStudy (anonymous):

radius= 1/2 diameter 2*radius= diameter 2*20=40 40=diameter

OpenStudy (anonymous):

i think its diameter is 40?